OpenStudy (anonymous):

Verify that the general solution satisfies the differential equation. Then find the particular solution that satisfies the initial condition. y = C1 sin(3x) + C2 cos(3x) y'' + 9y = 0 y = 7 when x = π/6 y' = 10 when x = π/6

OpenStudy (jango_in_dtown):

Hi do you have any problem in the first part ?

OpenStudy (anonymous):

No, what I'm missing & don't know how to find is C1 & C2

OpenStudy (jango_in_dtown):

show me the solution how far you reached

OpenStudy (anonymous):

i have 7 = C1 sin pi/6 + C2 cos pi/6 then 10 = C1 sin pi/6 + C2 cos pi/6

OpenStudy (jango_in_dtown):

its not pi/6.. it will be pi/2

OpenStudy (anonymous):

why

OpenStudy (jango_in_dtown):

3x=3pi/6=pi/2

OpenStudy (anonymous):

OH okay well I'm still unsure on how to solve

OpenStudy (jango_in_dtown):

Well you need to solve as follows

OpenStudy (jango_in_dtown):

see you made another mistake

OpenStudy (anonymous):

what would that be

OpenStudy (jango_in_dtown):

y = C1 sin(3x) + C2 cos(3x) gives y'=3C1 cos(3x)-3C2sin(3x) and so 10=3C1 cos pi/2 -3C2 si pi/2

OpenStudy (jango_in_dtown):

so we have the equations 7 = C1 sin pi/2 + C2 cos pi/2 =C1 and 10=-3C2 since sin pi/2=1 and cos pi/2=0 hence we have C1=7 and C2=-10/3

OpenStudy (jango_in_dtown):

@shelby.lane did you get the solution?

OpenStudy (anonymous):

I'm checking over it now

OpenStudy (anonymous):

yes! thank you!

OpenStudy (jango_in_dtown):

np.:)